Grrr i'm trying to figure out the mistake i made when i was solving this equation
How could I avoided it?
the equation 2(x-3) = x +3
First step in solving is 2x-2 =x +3

Answer by siouxie(6)

2(x-3) = x +3
First step in solving is 2x-2 =x +3
You didn't distribute the 2 in 2(x-3)
it should be 2*x -2*3 which is
2x-6 = x+3
Can you solve the rest? Get the variable on one side
2x-6 = x + 3
-x -x
------------
x -6 = 3
+6 +6
--------
x = 9
then subsitute 9 for x and check your answer